40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 341 of 357  Not logged in ELOG logo
ID Date Author Type Category Subject
  830   Tue Aug 12 21:38:19 2008 JohnUpdateLSCAccidental higher order mode resonances in the arms
Recently we had been having some trouble locking the full IFO in the spring configuration (SRC on +166).
It was thought that an accidental higher order mode resonance in the arms may have been causing problems.

I previously calculated the locations of the resonances using rough arm cavity parameters(Elog #690). Thanks to Koji
and Alberto I have been able to update this work with measured arm length and g factors for the y arm (Elog #801,#802).
I have also included the splitting of the modes caused by the astigmatic ETM. Code is attached.

I don't see any evidence of +166MHz resonances in the y arm.


In the attached plot different colours denote different frequencies +33, -33, +166, -166 & CR.
The numbers above each line are the mn of TEMmn.
Solid black line is the carrier resonance.
Attachment 1: HOMresonances.png
HOMresonances.png
Attachment 2: HOMarms2.m
%Check for accidental resonances of HOM in the arms (maybe due to
%sidebands). At the moment there is only data for the y arm.

clear all
close all
clc


%Stuff one might change often
modeorder = 0:5;          % Look for TEMmn modes where m,n run over modeorder
... 157 more lines ...
  829   Tue Aug 12 19:48:24 2008 JenneUpdateIOOPRM standoff is in....mostly
Yoichi, Jenne

The missing PRM standoff is now partially installed. The standoff is in, and the wire is in the groove, but we have not finished adjusting its position to make the PRM stand up straight. It turns out to be pretty tricky to get the position of the standoff just right.

We have set up a HeNe laser as an oplev on the flow bench (which we checked was level) in the clean assembly room, and are using it to check the pitch of the mirror. We set a QPD at the height of the laser, and are looking at the single-reflected light. When the single-reflected light is at the same height as the center of the QPD, then the mirror is correctly aligned in pitch. (Actually, right now we're just trying to get the single-reflected light to hit the diode at all...one step at a time here).

We'll continue trying to align the PRM's standoff in the morning.
  828   Tue Aug 12 12:21:13 2008 josephbConfigurationCamerasVariation in fit over 140 images for GC650 and GC750
Used matlab to calculate Gaussian fits on 145 GC650 images and 142 GC750 images. These were individual images (no averaging) looking at the PSL output from May 29th 2008. The GC650 and GC750 were looking at a split, but had different exposure values, slightly different distances to the nominal waist of the beam, and were not centered on the beam identically. Mostly this is a test of the fluctuations in the fit from image to image.

Note the mm refer to the size or position on the CCD or CMOS detector itself.
GC650

Mean
Amplitude  X center       Y center       X waist  Y waist  Background offset from zero
           position (mm)  position (mm)  (mm)     (mm)
0.3743     1.7378         2.6220         0.7901   0.8650  0.0047

Standard Deviation
Amplitude  X center       Y center       X waist  Y waist  Background offset from zero
           position (mm)  position (mm)  (mm)     (mm)
0.0024     0.0006         0.0005         0.0005   0.0003  0.00001

Std/Mean x100 (percent)
Amplitude  X center       Y center       X waist  Y waist  Background offset from zero
           position (mm)  position (mm)  (mm)     (mm)
0.6%       0.03%          0.02%          0.06%    0.04%    0.29%


GC750

Mean
Amplitude  X center       Y center       X waist  Y waist  Background offset from zero
           position (mm)  position (mm)  (mm)     (mm)
0.2024     2.5967         1.4458         0.8245   0.9194  0.0418

Standard Deviation
Amplitude  X center       Y center       X waist  Y waist  Background offset from zero
           position (mm)  position (mm)  (mm)     (mm)
0.0011     0.0005         0.0005         0.0003   0.0005  0.00003

Std/Mean x100 (percent)
Amplitude  X center       Y center       X waist  Y waist  Background offset from zero
           position (mm)  position (mm)  (mm)     (mm)
0.6%       0.02%          0.04%          0.04%    0.05%    0.07%
  827   Tue Aug 12 12:05:36 2008 YoichiUpdateComputersHP color printer is back
I restarted the HP printer server (a little box connected to the HP color laser) so that we can use the HP LaserJet 2550.
After this treatment, the printer spat out a bunch of pages from suspended jobs, many of these were black and white.
I think people should use the black-and-white printer for these kind of jobs, because the color printer is slow and troublesome.
  826   Mon Aug 11 19:09:28 2008 JenneDAQPEMSeismometer DAQ is being funny
While looking at the Ranger seismometer's output to figure out what our max typical ground motion is, Rana and I saw that the DAQ output is at a weird level. It looks like even though the input to the DAQ channel is being saturated, the channel isn't outputing as many counts as expected to Dataviewer.

Sharon and I checked that the output of the seismometer looks reasonable - sinusoidal when I tap on the seismometer, and the the output of the SR560 (preamp) is also fine, and not clipping. If I stomp on the floor, the output of the SR560 goes above 2V (to about 3V ish), so we should be saturating the DAQ, and getting the max number of counts out. However, as you can see in the first figure, taken when I was tapping the seismometer, the number of counts at saturation is well beneath 32768counts. (16 bit machine, so the +-2V of the DAQ should have a total range of 65536. +2V should correspond to +32768counts.) The second figure shows 40 days of seismometer data. It looks like we saturate the DAQ regularly.

I did a check of the DAQ using an HP6236B power supply. I sent in 1V, 2V and 2.2V (measuring the output of the power supply with a 'scope), and measured the number of counts output on the DAQ.

Input Voltage [V]Counts on DataviewerExpected counts from 16 bit machine
11898316384
22933132768
2.22934732768


I'm not sure why the +1V output more than the expected number of counts (unless I mis-measured the output from the power supply).

Moral of the story is...when the DAQ is saturated, it is not outputting the expected number of counts. To be explored further tomorrow...
Attachment 1: SeisDAQ.png
SeisDAQ.png
Attachment 2: SeisData.png
SeisData.png
  825   Mon Aug 11 15:07:49 2008 josephbConfigurationComputersProcyon aka fb40m switched to new switch
I've connected Procyon to the Prosafe 24 port switch with a new, labeled Cat6 cable. Quick tests with dataviewer shows that its working.
  824   Mon Aug 11 13:59:23 2008 josephbConfigurationComputers 
While poking around the crate, I noticed an error light on one of the c1susvme2 related boards was lit, while the corresponding light on the c1susvme1 was not. This confuses me as the c1susvme1 is the one having problems.

As a quick sanity check, I unplugged the ethernet connection from the c1susvme1 labeled board, and confirmed I couldn't log into it, and then plugged it back in, restarted it, and re-ran the startup script. This time c1susvme1 seemed to come up fine. Re-enabling the watchdogs doesn't seem to kick anything, and in fact seems to be bringing everything into line properly.

Although the error light on the c1susvme2 clk drvr board is still on. So I'm not sure what thats trying to tell us. Open to suggestions.
  823   Mon Aug 11 12:42:04 2008 josephbConfigurationComputersContinuing saga of c1susvme1
Coming back after lunch around 12:30pm, c1susvme1's status was again red. After switching off watchdogs, a reboot (ssh, su, reboot) and restarting startup.cmd, c1susvme1 is still reporting a max sync value (16384), occassionally dropping down to about 16377. The error light cycles between green and red as well.

At this point, I'm under the impression further reboots are not going to solve the problem.

Currently leaving the watchdogs associated with c1susvme1 off for the moment, at least until I get an idea of how to proceed.
  822   Mon Aug 11 11:36:11 2008 josephb, SteveConfigurationComputersc1susvme1 minor problems
Around 11 am c1susvme1 start having issues. Namely C1:SUS-PRM_FE_SYNC was railing at some large value like 16384 (2^14). I presume this means the computer was running catastophically late.

I turned off the BS and ITM watch dogs (the PRM was already off), tried hitting reset and sshing in, and running startup, but this didn't help. I then turned off the c1susvme2 associated watch dogs (MC1-3, SRM) and went out to do a hard reboot by switching the crate power off. c1susvme2 came back up fine, was restarted and associated watch dogs turned back on. However, c1susvme1 came back up without mounting /cvs/cds/.

As a test, I replaced the ethernet connection with a CAT6 cable to the Prosafe switch in 1Y6, and then ran reboot on c1susvme1. When it came back up, it had mounted properly, and I was able to run the ./startup.cmd file. At this point it seems to be happy. The new cable is in the trays coming in from the top of the 1Y4 and 1Y6 and approriately labeled.

Edit: Apparently ITMX and ITMY became excited after the reboot (perhaps I turned the watchdogs back on too early? Although that was after the DAQ light was listed as green for c1susvme). Steve noticed this when the alarms went off again (I had turned them off after the reboot seemed successful), and he damped them. Interestingly, the BS remained unexcited.
  821   Mon Aug 11 09:39:29 2008 ranaUpdatePEM2 years of temperature trend
Steve and I went around and inspected and then adjusted the thermostats and humidostats.

All the thermostats were set at 70F in 2005 by Steve. We adjusted the ones on the arms up to 72F
and set the one on the wall west of the framebuilder up to 74F (this area was consistently colder
than all the others and so we're over-correcting intentionally).
  820   Mon Aug 11 00:58:31 2008 ranaUpdatePEM2 years of temperature trend
The PSL RMTEMP alarmed again because it says the room temperature is 19.5 C. Steve said in
an earlier log entry that this is a false alarm but he didn't say why he thought so...

I say that either the calibration of the RMTEMP channel has drifted, the setpoint of the HVAC
has shifted, or there's a drift in the RMTEMP channel. I don't know what electronics exactly
are used for this channel so not sure if its susceptible to so much drift.

However, since the Dust Monitor (count_temp) shows a similar temperature decline in the
last two years I am inclined to blame the HVAC system.

The attached plot shows 2 years of hour-mean trend.
Attachment 1: Untitled.png
Untitled.png
  819   Sun Aug 10 16:57:02 2008 ranaSummaryPhotosPhotos from Vent 8/4 - 8/8
http://www.ligo.caltech.edu/~rana/40m/VentAug08/

I've added the D40 pictures from last week to this web page. I have done some cropping and
rotating to make things look better.

On page 3, there are some over head shots of the Michelson area so that one can use screw holes
to judge what the spacing between the suspensions is and also possibly the cavity lengths. Lets
also remember to measure the ITM-BS distance accurately using a tape measure or ruler while we
have the thing open.
  818   Fri Aug 8 17:54:52 2008 JenneUpdateSUSStandoffs and Guide Rods
After closer inspection of other small optics, it is clear that the guide rods should be above the standoffs on our small optics. Yoichi took a picture of the SRM that shows this clearly. This makes sense since the tension of the wire will make the standoff 'want' to go up during pre-epoxy adjustment, so the guide rod prevents the standoff from popping up and out.

Looking at the side of the PRM without the groove, it looks like there is lots of space between the guide rod and the alignment etch in the glass, so we can just place a standoff directly under the guide rod that is present.

A spare standoff is being shipped tomorrow morning, so we should have it by Monday for installation on the PRM.
Attachment 1: SRM_Standoff_and_guide.JPG
SRM_Standoff_and_guide.JPG
  817   Fri Aug 8 15:10:35 2008 YoichiUpdateSUSNo groove in the stand-off ... wait, it is not even a stand-off !
I tried to find the missing stand-off and the guide rod in the BS chamber, but I couldn't.
  816   Fri Aug 8 13:29:54 2008 YoichiUpdateSUSNo groove in the stand-off ... wait, it is not even a stand-off !
Yoichi, Steve, Seiji

We took magnified pictures of the stand-offs of the PRM.
Attm1: North side stand-off.
Attm2: South side stand-off.
Attm3: Zipped file of the full pictures.

We found no groove in the south side stand-off.
After some discussion, we concluded that it is actually a guide rod. You can see it from the size difference (the magnification is the same for the two pictures).
The stand off on the south side is missing (fell off, ran away, evaporated or whatever ...).
Also we noticed that the guide rod on the north side is missing.

We have to find a stand-off and place it on the south side.
Seiji suggested that it is better to put a guide rod next to the north side stand-off, otherwise the stand-off itself is too weak to hold the load.
He also said that the PRM was installed after he left, so it was not his fault.
Attachment 1: north-standoff-preview.jpg
north-standoff-preview.jpg
Attachment 2: south-standoff-preview.jpg
south-standoff-preview.jpg
Attachment 3: No-groove.zip
  815   Fri Aug 8 12:21:57 2008 josephbConfigurationComputersSwitched X end ethernet connections over to new switch
In 1X4, I've switched the ethernet connections from c1iscex and c1auxex over to the new Prosafe 24 port switches. They also use the new cat6 cables, and are labeled.

At the moment, everything seems to be working as normally as it was before. In addition:

I can telnet into c1auxex (and can do the same to c1auxey which I didn't touch).
I can't telnet into c1iscex (but I couldn't do that before, nor can I telnet into c1iscey either, and I think these are computers which once running don't let you in).
  814   Fri Aug 8 11:04:34 2008 SharonUpdate MCL Wiener filter
I took some old data from Rana and converted the units of the Weiner filter to m/m so to make the effect of the seismometer and accelerometers more obvious.

The data is in counts, and so to convert to m this is what I did:

%%% MC_L calibration
v_per_counts = 5/32768;
v_per_v = 3;
amp_per_N = 1/0.064;

%%% Accelerometers calibration
v_per_counts_acc = 61.045e-6;
g_per_v = 9.8/100;

%%% Seismometer calibration
v_per_counts_seis = 61.045e-6;
m_per_s_per_s_per_volt = 9.8/100;
m_per_v_per_s = 1/345;



for jj=1:6
hfmatm(:,jj)=hfmat(:,jj).*(v_per_counts.*v_per_v.*amp_per_N.*f)./(v_per_counts_acc*g_per_v); %%% accelerometers' data
end
hfmatm(:,7)=hfmat(:,7).*(v_per_counts.*v_per_v.*amp_per_N)./(v_per_counts_seis*m_per_v_per_s); %%% Seismometer data
Attachment 1: m_per_m.png
m_per_m.png
  813   Fri Aug 8 10:58:05 2008 josephbConfigurationCamerasCameras and gstreamer
In regards to camera failure:
1) I forgot to reconnect that particular camera to the network (my fault) so thats why it was failing.

2) Even with the correct camera connected, I've realized at full frame rate, op440m is going to get a few frames and then fail, as I don't think it has a fast enough ethernet card. It will work on Rosalba, and will also work ssh-ing from Rosalba because it is using a new ethernet card. It also works on my laptop, which is where I originally tested the command. One way to get around this is to increase the time between pictures, by changing -l 0 to -l 1 (or higher), where the number after the "ell" is the number of seconds to wait between frame captures.

3) What I should do is figure out the UDP transmission plugins for gstreamers and compress first (using the theoraenc since it gets compression ratios of better than 100:1) and transmit that over the network.

I have since reconnected the camera, so it should work on Rosalba and any sufficiently well connected computer. For other machines like linux2 or op440, try the following line:

Running the following command on Mafalda (via ssh -X mafalda) or Rosalba while in /cvs/cds/caltech/target/Prosilica/40mCode/SnapCode/

CamSnap -F 'Mono8' -c 44058 -E 10000 -X 0 -Y 0 -H 480 -W 752 -l 1 -m 100 | gst-launch-0.10 fdsrc fd=0 blocksize=360960 ! video/x-raw-gray, height=480, width=752, bpp=8,depth=8,framerate=1/1 ! ffmpegcolorspace ! ximagesink

This will be at a much slower frame rate (1 per second) but should work on any of the machines. (Tested on linux2).
  812   Fri Aug 8 09:54:10 2008 ranaUpdateCamerasNew code + gstreamer allows for easy saving and compression of images

Quote:
Modified the CamSnap code to output the image data stream to standard out. This can then be piped into a gstreamer plugin and then be used


Didn't work; Prosilica has only 1 "l". Even so, sshing from op440m to mafalda, I got this:
mafalda:SnapCode>CamSnap -F 'Mono8' -c 44058 -E 5000 -X 0 -Y 0 -H 480 -W 752 -l 0 -m 300 | gst-launch-0.10 fdsrc fd=0 blocksize=360960 ! video/x-raw
-gray, height=480, width=752, bpp=8,depth=8,framerate=30/1 ! ffmpegcolorspace ! theoraenc ! oggmux ! filesink location="./testVideo.ogm"
Setting pipeline to PAUSED ...
Pipeline is PREROLLING ...

** (gst-launch-0.10:27121): WARNING **: Size 60 is not a multiple of unit size 360960
Caught SIGSEGV accessing address 0x487c
ERROR: from element /pipeline0/ffmpegcsp0: subclass did not specify output size
Additional debug info:
gstbasetransform.c(1495): gst_base_transform_handle_buffer (): /pipeline0/ffmpegcsp0:
subclass did not specify output size
ERROR: pipeline doesn't want to preroll.
Setting pipeline to NULL ...
#0  0xffffe410 in __kernel_vsyscall ()
#1  0xb7deddae in __lll_mutex_lock_wait ()
#2  0xb7de9aac in _L_mutex_lock_51 () from /lib/tls/i686/cmov/libpthread.so.0
#3  0xb7de949d in pthread_mutex_lock ()
#4  0xb7e452e0 in g_static_rec_mutex_lock () from /usr/lib/libglib-2.0.so.0
#5  0xb7f1fa08 in ?? () from /usr/lib/libgstreamer-0.10.so.0
#6  0x080c1220 in ?? ()
#7  0x00000001 in ?? ()
#8  0x0809586c in ?? ()
#9  0x00000001 in ?? ()
#10 0x08095868 in ?? ()
#11 0xb7f7a2a8 in ?? () from /usr/lib/libgstreamer-0.10.so.0
#12 0xb7e8da80 in ?? () from /usr/lib/libglib-2.0.so.0
#13 0xb7f7a2a8 in ?? () from /usr/lib/libgstreamer-0.10.so.0
#14 0xb7f7a2a8 in ?? () from /usr/lib/libgstreamer-0.10.so.0
#15 0x00000000 in ?? ()
Spinning.  Please run 'gdb gst-launch 27121' to continue debugging, Ctrl-C to quit, or Ctrl-\ to dump core.
Caught interrupt -- 
  811   Thu Aug 7 17:32:23 2008 JenneUpdateSUSAfternoon PRM activities
Rana, Jenne, Yoichi, Dmass

After Yoichi confirmed this morning that the wire was in both grooves, Rana attempted to lift the PRM a tiny bit, and twist it around (very gently of course) to see if we could make the wire slip back to its nominal position underneath the optic. On the first attempt, the wire ended up slipping the wrong way, causing slightly more tilt. On another attempt, the wire came out of the groove nearest the chamber door by about 0.5mm. We got the wire back in the groove by slightly lifting the optic, and pushing the wire back in. Then, on further attempts at making the wire slip back to its nominal position, the wire came out of the groove farthest from the chamber door. It is very difficult to get at that side of the PRM, because the table is crowded, and it is on the far side of the optical table from the chamber door. We decided to pull the PRM out of the chamber. Rana clamped the mirror into its cage using the earthquake stops and removed the OSEMS, and then we pulled the mirror out. We put it on a cart that was covered with foil and had a little foil house for the optic cage. We rolled the mirror+cage over to the flow bench at the end of the y-arm.

We saw that the wire is no longer even on the standoff (~3mm away from the groove) on the side that was farthest from the chamber door.

Since we have not confirmed that we have spare wire and spare magnets (and due to the time of day), we have decided to cover the cage with some foil, while it is sitting on the flow bench, and we'll fix the wire in the morning.
  810   Thu Aug 7 12:20:52 2008 YoichiUpdateSUSPRM stand-offs and wire
We removed the side OSEM of the PRM so that we can see the stand-off on the farther side.

Attachment 1: Farther side stand-off from an angle before removing the OSEM
Attachment 2: Farther side stand-off through the empty OSEM hole.
Attachment 3: Near side stand-off

The wire is definitely in the near side stand-off groove.
Probably the wire is in the groove also on the farther side.
Attachment 1: IMG_1456.JPG
IMG_1456.JPG
Attachment 2: IMG_1478.JPG
IMG_1478.JPG
Attachment 3: IMG_1470.JPG
IMG_1470.JPG
  809   Thu Aug 7 11:54:26 2008 josephbConfigurationCamerasNew code + gstreamer allows for easy saving and compression of images
Modified the CamSnap code to output the image data stream to standard out. This can then be piped into a gstreamer plugin and then be used to save, encode, transmit, receive, slice, dice and or mangle video (or virtually any type of data stream).

The gstreamer webpage can be found at: http://www.gstreamer.net/

Under documentation you can find a list off all available plug-ins. Some good, some bad, some ugly.

Running the following command on Mafalda (via ssh -X mafalda) or Rosalba while in /cvs/cds/caltech/target/Prosilica/40mCode/SnapCode/

CamSnap -F 'Mono8' -c 44058 -E 15000 -X 0 -Y 0 -H 480 -W 752 -l 0 -m 1000 | gst-launch-0.10 fdsrc fd=0 blocksize=360960 ! video/x-raw-gray, height=480, width=752, bpp=8,depth=8,framerate=1/1 ! ffmpegcolorspace ! ximagesink

This command will create a window which displays what the camera with UID 44058 is looking at. It will display 1000 images, then quit. (You can switch the -m 100 to -i to just have it continue until the process is stopped).

You can also encode the data into compressed format and save it in a media file. The following command line will encode the images into an ogg media file (.ogm), which can be played with the totem viewer (available on Rosalba or almost any machine running Ubuntu or Centos) or any other viewer capable of handling ogm files. By switching the plugins you can generate other formats as well.

The compression is good, putting 300 images normally about 500K individually uncompressed to about 580K as a single file.

The following command line was used to generate the attached video file:

CamSnap -F 'Mono8' -c 44058 -E 5000 -X 0 -Y 0 -H 480 -W 752 -l 0 -m 300 | gst-launch-0.10 fdsrc fd=0 blocksize=360960 ! video/x-raw-gray, height=480, width=752, bpp=8,depth=8,framerate=30/1 ! ffmpegcolorspace ! theoraenc ! oggmux ! filesink location="./testVideo.ogm"

Currently looking into plugins which allow you to pull individual frames out of a video file and display or save them in a variety of formats. This would allow us to save long term images in compressed video format, and then pull out individual frames as needed.

Also need to look into how to "T" the streams, so one can be displaying while another encodes and saves.
Attachment 1: testVideo.ogm
  808   Thu Aug 7 10:27:59 2008 ranaUpdateSUSFree swinging OSEM spectra
Sometimes we see extra peaks in the OSEM spectra coming from a beat between the regular eigenmodes.
This probably comes from the OSEM shadow sensor not being entirely linear - the nonlinearity is
greatly increased if the magnet is not perfectly centered in the LED beam. So the beats are
probably there at some level in all of them; usually below the noise.
  807   Thu Aug 7 10:07:13 2008 YoichiUpdateSUSFree swinging OSEM spectra
Looks like there are more extra peaks in the SRM than other optics.
Maybe because it is closer to the door ?
Attachment 1: FreeSwingSpectra.pdf
FreeSwingSpectra.pdf FreeSwingSpectra.pdf FreeSwingSpectra.pdf FreeSwingSpectra.pdf FreeSwingSpectra.pdf FreeSwingSpectra.pdf FreeSwingSpectra.pdf FreeSwingSpectra.pdf
  806   Wed Aug 6 22:19:07 2008 YoichiUpdateSUSBS alignment
Koji, Yoichi

We realized that we did not pay attention to the BS alignment while working on the alignment of the ITMX today. Because we were injecting the ALM laser (absolute length measurement laser) from the AS port, the ITMX alignment depends on the BS alignment.
The BS optical lever was not centered and the sum was about 2000cnt, which is low compared, for example, to the SRM oplev.
So we were not sure if the BS was in a good alignment or not.
So we decided to move the BS to center the QPD.
In doing so, we also moved the ITMX so that we do not lose the ALM laser beam coming back to the AS port.
When the BS oplev was centered, the sum of the QPD was still about 2000. So it was not far off centered.
After the tweaking, we were able to see some interference between the light reflected by the ITMY and ITMX at the AS port (actually this is the bright port for the ALM laser). By tweaking the ITMY, we were able to see Michelson fringes at the AS port.
If we believe the ALM laser alignment is still good after the vent, the ITMX, ITMY, BS and SRM should be now in a good alignment condition.
The OSEM values for the ITMX, BS, SRM seem to be ok (0.9+/-0.2). The ITMY LL is a bit low (~ 0.45).
  805   Wed Aug 6 19:01:15 2008 AlbertoUpdateGeneralITMX and SRM OSEM post-earthquake diagnostic
Koji, Yoichi, Alberto

Today we reset the OSEMs on ITMX and SRM in order to be centered when the mirrors are aligned to the IFO beam. Since the PRM is still out of order, we used the beam from NPRO laser of the absolute length measurement experiment as it is injected through the AS port.
That’s how we did it:

1) We aligned the SRM so that the reflected beam from the NPRO was at the camera after at the AS port.

2) We traded off the alignment of SRM in order for the reflected beam at the camera to have a nice shape, avoiding any clipping from the optics, and for the optical lever to be not too far from zero. The final alignment for SRM, as read on the sliders on the MDM screen, is: Pitch=1.1650, Yaw=1.4674.

3) We aligned ITMX checking out by an IR card that the incoming and the reflected main beam in between ITMX and the BS matched. The alignment of the two beams was improved checking the matching after the SRM. The final alignment for ITMX, as read on the sliders on the MDM screen, is: Pitch=-1.2937, Yaw=-0.9890.

4) After the alignment of SRM and ITMX these were the voltages at the OSEMs:

SRM
UL=0.957
UR=1.254
LR=0.768
LL=0.620
Side=0.958

ITMX
UL=1.144
UR=1.360
LR=0.591
LL=0.325
Side=-----

5) Finally we centered the OSEMs on both mirrors and we read these voltages:

SRM
UL=0.939
UR=0.994
LR=0.782
LL=0.938
Side=0.953

ITMX
UL=0.918
UR=0.891
LR=0.887
LL=0.875
Side=0.883
  804   Wed Aug 6 13:57:44 2008 MashaSummaryAuxiliary lockingweekly summary
Finished second progress report.

Working on improving the sensitivity of the Mach Zehnder to more accurately measure the fiber noise. Making more stable mounts that have fewer degrees of freedom/springs and are more solid should get rid of their vibrational modes and help with the noise. I found a good mount for the Faraday Isolator, and John and Aidan helped me to make the solid aluminum blocks to mount the fiber couplers. I'll also replace the laser feet with a similar solid mount. I will also get a plastic box to block out acoustic noise/air currents/etc. Am starting to couple into the fiber again; now I am using polarization maintaining fiber.

I am starting to plan the fiber noise cancellation setup, and thinking about the noise sources and their effects.
  803   Wed Aug 6 13:15:57 2008 YoichiUpdateSUSSRM ETMX freeswing spectra
After yesterday's work on the SRM, I took free swinging spectra of SRM.
The eigen modes look ok. But there are many other peaks which were not present in vacuum.
Some of those peaks may be resonances of the air inside the chambers and the pipes.
However, the peaks around 0.2Hz are too low frequency for those air compression modes.
I took the ETMX spectra at roughly the same time. I chose ETMX because we have not touched it after the vent.
ETMX also shows some extra peaks but the frequencies are different.
Attachment 1: SRM-ETMX-freeswing.pdf
SRM-ETMX-freeswing.pdf SRM-ETMX-freeswing.pdf
  802   Wed Aug 6 11:43:52 2008 KojiUpdateGeneralAbs. Len. Meas. ~ analysis of the TEM01 scan
Analysis of the data on August 3th ~ Part 2

o I already have reported that the resonant freq of TEM10 and TEM01 split.

o Again, note that TEM10/01 were arranged almost in the horizontal/vertical by the observation of the video.

o The peaks of TEM10 and TEM01 were fitted with the same method as of TEM00.

o The peak freqs were:
f_TEM10: 5087040 Hz +/- 20 Hz
f_TEM01: 5068322 Hz +/- 15 Hz
The split is 18.7kHz.

o The additional parameter from the previous entry:
f_TEM00: 3879252 Hz +/- 9 Hz
L_yarm: 38.6462 m +/- 0.0003 m

o Radius of curvature
Rx = L /(1-Cos^2(Pi (f_TEM10 - f_TEM00) / (c/L/2) ))
Ry = L /(1-Cos^2(Pi (f_TEM01 - f_TEM00) / (c/L/2) ))

from these formula we get the value
Rx = 56.1620 +/- 0.0013 [m]
Ry = 57.3395 +/- 0.0011 [m]
Attachment 1: TEM01fit.png
TEM01fit.png
  801   Wed Aug 6 11:10:34 2008 KojiUpdateGeneralAbs. Len. Meas. ~ analysis of the TEM00 scan
Analysis of the data on August 3th ~ Part 1

From the measurement of the 5 FSRs, the FSR frequency for the Yarm cavity was estimated as
f_FSR = 3878678 Hz +/- 30 Hz
and the Yarm length is
L_yarm = 38.6462 m +/- 0.0003 m
This is the precision of 8ppm. In my opinion, this is a satisfactory result for our purpose.
Y-arm length
e-log    length [m]
-----------------------------
 556(2008-Jun-24)  38.70    +/- 0.08    Cavity swinging measurement
 556(2008-Jun-24)  38.67    +/- 0.03    Tape & photo
 776(2008-Jul-31)  38.640   +/- 0.007   Beam injection, poor PLL, Transmitted DC
 782(2008-Aug-02)  38.6455  +/- 0.0012  Beam injection, independent PLL, Transmitted DC
 787(2008-Aug-04)  38.64575 +/- 0.00037 Beam injection, independent PLL, Transmitted RF
this(2008-Aug-04)  38.6462  +/- 0.0003  Beam injection, independent PLL, Transmitted RF, five FSRs, freq calibrated
-----------------------------
----------------
o According to the entry 795, all of the scan frequency was calibrated.
o The five peaks of the scanned data for TEM00 were fitted. Each peak was fitted by the following formula:

V(f) = A / Sqrt(1 - ((f-f0)/fc)^2)

Variable
f: scan frequency

Parameters
A: peak amplitude
f0: center frequency
fc: half bandwidth of the peak for -3dB

o The results are shown in the attached figure 1. They look very similar each other but they are different plot! The fittings were extremely good. The center frequencies estimated were as follows:
FSR1:  3879251.9 Hz +/-  8.8 Hz
FSR2:  7757968.1 Hz +/- 10.8 Hz
FSR3: 11636612.9 Hz +/- 10.2 Hz
FSR4: 15515308.1 Hz +/-  8.7 Hz
FSR5: 19393968.7 Hz +/-  8.4 Hz
o The FSR frequencies were fitted by a line. The fitting and the residuals are shown in the attached figure 2.
The fitting results were

f_FSR(n) = 586.4 + 3878678 * n

This means that:
o FSR frequency was 3878678 [Hz].
o The lock of the carrier had detuning of 586 [Hz].

The detuning of the carrier from the resonance can be explained by the alignment drift. In deed, at the end of the measurement, decrease of the transmitted power by -15% was found. Then, the frequency of the 1st FSR was measured before and after the alignment adjustment. This changed the frequency of the FSR1 by 350Hz. This change could not be explained by the cavity length change as this is too big (~3.5mm).

Actually, the spacing of the cavity length is more stable. The residual is rather scattered with in 20-30Hz. So, I took the error of 30Hz as the whole precision of the frequency measurement that includes the fluctuation of the alignment, the cavity length itself, and so on. This yields the FSR and the cavity length of
f_FSR = 3878678 Hz +/- 30 Hz
L_yarm = 38.6462 m +/- 0.0003 m .
Attachment 1: TEM00fit.png
TEM00fit.png
Attachment 2: TEM00FSRfit.png
TEM00FSRfit.png
  800   Tue Aug 5 17:56:23 2008 AlbertoConfigurationGeneralSRM and PRM inspection
Yoichi, Koji, Rana, Steve, Alberto

Today we opened the BSC to inspect the optics, and in particular the SRM and PRM.
We found that one of the side magnets of the SRM was broken and a piece of it fell and got stuck to the LR magnet.
We removed the LR OSEM and took off the broken part with tweezers. Since we couldn’t replace the magnet on the side,
we decided to just switch the OSEM to the other side were a second magnet was available. Then we centered the OSEMs.
Using the optical levers we aligned both the ITMX and the SRM so that now we have to center again the OSEMs on both.

The PRM was visibly tilted and it was out of the range of the OSEMs. To try to fix the tilt we lift it up a little
with the screws on the bottom and pushed it with the third screw on top. That had the effect of making the mirror
tilt to the opposite direction. We looked at the wires (see attached picture) and it seemed centered on the side
of the mirror.

Tomorrow we are going to reset the OSEMs on ITMX and SRM and then we’re going to try to fix the tilt on PRM.
Attachment 1: IMG_1434.JPG
IMG_1434.JPG
Attachment 2: IMG_1456.JPG
IMG_1456.JPG
  799   Tue Aug 5 12:52:28 2008 YoichiUpdateSUSITMX, SRM OSEM spectra
Free swinging spectra of ITMX and SRM.
ITMX seems to be ok after yesterday's work, though the OSEM DC values are still a bit off from the normal value of 0.9.
(ITMX OSEM values: UL=1.12, UR=1.38, LR=0.66, LL=0.41, SIDE=0.66)
SRM is still clearly wrong.
Attachment 1: ITMX-2008_08_05-morning.pdf
ITMX-2008_08_05-morning.pdf
Attachment 2: SRM-2008_08_05-morning.pdf
SRM-2008_08_05-morning.pdf
  798   Tue Aug 5 10:56:05 2008 AlbertoConfigurationGeneralITMX chamber opened and mirror released
D-Mass, Steve, Rana, Koji, Yoichi, Alberto,
We opened the ITMX chamber to check the optics after last week earthquake. In particular, from the spectra, ITMX seemed to be stuck and had to be released again. When we inspected the mirror, we found that it wasn’t necessary to touch it. It had become free again during the vent thanks to the change of conductivity in the air inside during the vent.
We checked the magnets and they seemed to be fine.
A couple of stop screws had lost the rubber on their tips, although we don’t know if that was due to the earthquake.
We also took advantage of the opening to center the LR and the left OSEMs in the mirror to their zero.
Inspecting the table we found a couple of things not totally clear on the configuration of the optics in the table. In particular we found a beam dump located too close to the ifo beam. Eventually we found out that the dump was meant to block a ghost beam coming from the ITM. A better location should probably be figured out for that. We also found that the POXM1 mirror designed to have the maximum reflectivity for the P polarization of the beam at 45 degrees is mounted so that the incident beam is at 22 degrees. This cause the beam to be 90% transmitted and only 10 percent reflected to POX. The transmitted beam appears at ther BSC chamber.

The ifo beam passes so close to the POXM1 mirror so that it can be clipped by its large metal frame ring. The beam at that point is about 6mm large and the ring is about 1cm thick so that we could gain some distance with a different mount.
  797   Tue Aug 5 10:23:00 2008 steveUpdateSUSearthquake and venting effects
atm 1, EQ
atm 2, vent 7 days later: venting kicks optic into place to be free,
PRM: LR magnet gets pushed in and it is stocked, side in free
Attachment 1: eq4h.jpg
eq4h.jpg
Attachment 2: vent4hr.jpg
vent4hr.jpg
  796   Tue Aug 5 02:39:55 2008 KojiConfigurationGeneralAbs. Len. Meas. ~ Optical Layout on the AP / PSL table 2008-Aug-05
Here are the PDF and the PNG of the AP and PSL table layouts.
After this photo, the squeezing setup at the AP table was removed.
Attachment 1: optical_layout_ap_table3.png
optical_layout_ap_table3.png
Attachment 2: optical_layout_ap_table3.pdf
optical_layout_ap_table3.pdf
Attachment 3: optical_layout_PSL_table1.png
optical_layout_PSL_table1.png
Attachment 4: optical_layout_PSL_table1.pdf
optical_layout_PSL_table1.pdf
  795   Tue Aug 5 00:05:57 2008 KojiUpdateGeneralAbs. Len. Meas. ~ IFR2023A calibration
Work log on August 4th

o IFR2023A (Marconi) was calibrated by the SR620 frequency counter which is locked to the GPS signal.
o The frequency of the IFR2023A was scanned from 1MHz to 20MHz with 1MHz interval. The readout of the frequency counter was recorded.
o The linear fit was taken.
f_freq_count = K0 + K1 * f_IFR [Hz]
K0 = 0.00        +/- 0.02
K1 = 0.999999470 +/- 0.000000001

o So, the IFR seems to have -0.5ppm systematic error.
  794   Mon Aug 4 22:31:10 2008 KojiUpdateGeneralAbs. Len. Meas. ~ Simple Test for TEM01/10 split
Work log on August 3rd - Part3

Question:
o The TEM01 and TEM10 of the Yarm were found to split with 19kHz separation. Is this true?
o In which direction the eigenmodes are?

Thought:
o The separation of 19kHz is a kind of too big because the cavity bandwidth is several kHz.
o This means that "TEM01 and TEM10 can not resonate at the same time (by the PSL beam)".

Test:
o Imagine we are just using the PSL beam and playing with an arm cavity.
o Tilt the end mirror in pitch. Resonate the TEM01 mode (8-shaped).
o Then tilt the end mirror in yaw.
o a) If the resonances are degenerated within the bandwidth of the laser, it rotates freely.
o b) If the resonances splits, the tilt in yaw does not change the shape. Then suddenly jumps to TEM10 (by an accident).

Result:
o The shape does not change. Just jumps to the other mode. (The case above b.)
o The eigenmode looked like quite horizontal and vertical.

Conclusion: the mode really splits.
Attachment 1: TEM01_10.png
TEM01_10.png
  793   Mon Aug 4 21:48:24 2008 KojiUpdateGeneralAbs. Len. Meas. ~ Scan for TEM00/01/10
Work log on August 3rd - Part2

o I tried to measure the frequency of the FSRs using TEM00 resonances. Also search of TEM01/TEM10 resonances were tried.

-----------
Measurement for TEM00

o The frequency of the injection beam was scanned from 2MHz to 20MHz using the LabVIEW panel with GPIB. The 1st figure attached below is the result of the scan. Equispaced peaks were found as expected. The interval of the peaks are about 3.89MHz. Each peaks were measured with freq intervals of down to 50Hz. I will analyze the center frequency of the peaks precisely later in order to have a final result.

Measurements for TEM01/TEM10

o The beam injection technique is thought to be useful for measureing the frequency of the higher-order resonances. In order to measure the higher-order resonances the modifications of the experimental setup were applied as below.

1) For TEM10 (the beam like "OO" shape), a razor blade which blocked the horizontal half of the transmitted beam was placed. We needed to disturb half of the beam because the beat between the PSL TEM00 and the injection TEM01 cancels if the PD receives all of the light.

2) The injection beam is slightly misaligned in the horizontal direction in order to enhance the coupling of the injection beam to the cavity TEM01 mode.

3) For TEM01 (the beam like "8" shape), a razor blade cutting the vertical half and the misalignment of the inj beam in the vertical direction are applied.

o The frequency of the injection beam was scaned from 1st FSR of TEM00 in the upward direction. The alignment of the arm cavity was left untouched during the measurement. As shown in the 2nd figure attached below, the resonances were found about 1.19MHz away from the TEM00, but they are separated by about 19kHz(!). This could be split of the degenerated modes which corresponds to the difference of the mirror curvature in two directions! This difference is something like 56 m and 57 m. Can you believe this?

(To be continued to the next entry)
Attachment 1: TEM00.png
TEM00.png
Attachment 2: TEM01.png
TEM01.png
Attachment 3: knife_edge.png
knife_edge.png
  792   Mon Aug 4 16:20:20 2008 DmassConfigurationPhotosITMX magnet position relative to OSEMS
We have vented, and taken the following pics of the magnets to document their position before we ruin everything.
Attachment 1: DSC_0151.JPG
DSC_0151.JPG
Attachment 2: DSC_0150.JPG
DSC_0150.JPG
  791   Mon Aug 4 13:43:02 2008 YoichiSummaryPSLFSS loop calibration
As a part of the effort to repair the FSS loop bandwidth, I tried to calibrate the FSS loop.

First, I scanned the MOPA frequency by injecting a triangular wave into the ramp-in of the FSS box, which goes to the PZT of the NPRO.
The first attachment shows the transmitted light curve (pink one) along with the PDH signal (light blue).
The sweep was very slow (0.1Hz for 2Vp-p). From this measurement, the FWHM was 6.8e-3V. Then fpol = FWHM/2=3.4e-3V, where fpol is the cavity pole frequency.
So the PZT's DC response is 294*fpol/V. If we use the canonical fpol=38kHz, it is 11.172MHz/V.

Then I tried to measure the cavity pole. First I tried the cavity ring down measurement, by blocking the beam abruptly. Unfortunately, my hand was not fast enough.
The ring down shape was not an exponential decay.
I then locked the reference cavity only using the PZT with very narrow bandwidth (UGF=2kHz). I injected signal into the external modulation input of the 80MHz VCO
for the AOM. The second attachment shows the transfer function from this input to the IN2 (mixer output monitor port) of the FSS servo box.
To plot this, I corrected the measurement for the open loop TF (i.e. multiplying the measured TF with (1+G)), and other filters in the path (8MHz LPF after the ext. mod.
input of the 80MHz VCO, and an RCL network after the mixter). The gain looks like a cavity pole, but the phase decreases very rapidly.
If you look at the third attachment showing a wider band transfer function, there are notches at 1.8MHz and above. I couldn't find this kind of filter in the schematic.
Maybe this is the RFPD's bandpass filter. I will check this later. From these plots, it is difficult to tell the cavity pole frequency. From the -3dB point, fpol is around 83kHz,
but from the phase=-45deg point, fpol is around 40kHz.

Finally, I calibrated the cavity's optical gain by locking the Ref. Cavity with only PZT, and injecting a signal into the loop.
The signal was injected from Test-In2 of the FSS servo box and the transfer function from the PZT output signal (TP10) to IN1 (mixer output) was measured.
The transfer function was corrected for a 10Hz LPF after TP10.
The attachment4 shows a nice flat response up to 30kHz. Above 30kHz, the measurement is too noisy. The optical gain at DC is about 22dB from the PZT drive to the error signal (IN1).
Using fpol=38kHz, it means 887kHz/V calibration factor for the signal at IN1. There is a mixer output monitor DAQ channel in the FSS but it seems to be not working at the
moment. I will look into this later. There is a gain of 10dB between IN1 and the mixer monitor channel.
By looking at the phase response of the attachement4, there is a cavity pole like behavior around 30kHz. If we assume the PZT response is flat up to this frequency, it is
roughly consistent with fpol=38kHz.

I was not able to take a sensible spectrum of IN1 using the network analyzer. When the FSS servo was engaged, the signal was too small.
I will try to use an AF spectrum analyzer later to get a calibrated spectrum.
Attachment 1: P7310048.JPG
P7310048.JPG
Attachment 2: cavity-response.pdf
cavity-response.pdf
Attachment 3: cavity-response2.pdf
cavity-response2.pdf
Attachment 4: cavity-gain.pdf
cavity-gain.pdf
  790   Mon Aug 4 12:12:20 2008 steveConfigurationVACthe ifo is at atm
The 40m vac envelope was vented this morning.
P1 is at 760 Torr
Attachment 1: vent3h.jpg
vent3h.jpg
Attachment 2: Untitled.png
Untitled.png
  789   Mon Aug 4 05:23:57 2008 KojiUpdateGeneralAbs. Len. Meas. ~ Measurement for Y-arm completed
Finally, I have completed the abs length and g-factor measurements for Y-arm.
>>>GO FOR THE VENT<<<

I will report the results later.

Some notes on the status:
o Y-arm was aligned at the end of the experiment by the script. The values were saved.

o At the AP table, the injection beam and the flipper were left aligned so that the inj. beam can be used as a reference of the SRM and the ITMs. But the shutter of the NPRO was closed.

o The experiment setup was mostly left at the side of the AP table. I tried not to disturb the walk as much as possible.

o The long cable from the Y-end was wound and placed at the Y-end. The knife-edge was left on the Y-end bench. It is not disturbing any beam.
  788   Mon Aug 4 00:56:07 2008 KojiHowToGeneralAbs. Len. Meas. ~ Auto freq scanner with GPIB
Work log on August 3rd - Part1

o Yesterday I was too much tired of changing the RF frequency, reading peaks on the RF spectrum, and writing the values. Rana saw me and thought I was such poor that he gave me an USB-GPIB adapter.

o I dig into the internet for the manuals of the adapter, IFR2023A(Marconi), and HP8591E(RF spectrum analyzer) in order to learn how to use them.

o I had LabVIEW installed on my laptop. Finally I understand how to use that adapter (by Agilent) with LabVIEW. I made a small program to scan the frequency of IFR2023A, and read the peak values from HP8591E. It is unfortunate that there is no LabVIEW in the 40m lab. I think I can make an independent executable which does not need the LabVIEW itself. Give me some time to understand how to do it.
Attachment 1: freq_scan.png
freq_scan.png
  787   Mon Aug 4 00:37:58 2008 KojiUpdateGeneralAbs. Len. Meas. ~ RF PD at the Y end / Manual frequency scan
Work log on August 2nd

o Just remind you:
The idea of the absolute length measurement was to detect an RF beat between the injection beam and the PSL beam by resonating both of the beams to the cavity at the same time, but on different londitudinal modes. From the frequency separation between the two beams, we get the FSR of the cavity. In order to have an injection beam with stable frequency separation, a heterodne interferometer was built at the PSL table, and the PLL servo is used to control and stabilize the frequency of the inj. beam.

----------

o An RF PD (Tholab PDA255) and a steering mirror were placed at the Yarm END. Fortunately, I found that an unused BS was already in the optical path. There was a beam block which dump the reflection of the BS and some stray lights of the OPLEV. I moved the beam block to make the BS reflection available, as well as to block the OPLEV stray light still (Photo1). In order to have the RF signal from the PD, a long BNC cable was laid along the Yarm. I did't know any better idea than this. Don't blame me.

o To have an intuitive interpretation of the beat frequency, the injection beam was set to be at higher frequency than the PSL beam. How did I confirm this? When the crystal temp (LT) of the NPRO was tuned to be higher, the beat frequency got lower.

o Frequency of the PLL was manually swept at around 15.51MHz where the 4th FSR was expected to be found. I could see strong RF peak at that frequency! When I tuned the PLL frequency, the peak height changes dramatically! Too cool!

o The amplitude of the RF peak was measured by an RF spectrum analyzer. I did all of this scan by my hands and eyes. The center frequency of the 4th FSR was 15.5149MHz. From the eye I would say the error is +/-150Hz. It is OK so far although I am not sure statistically this is correct or not. This corresponds to the length of 38.64575 +/- 0.00037 [m].

o All of the past measurements are fairly consistent.
Y-arm length
e-log              length [m]           Measurement Conditions
----------------------------------------------------------------------------------------
 556(2008-Jun-24)  38.67    +/- 0.03    Cavity swinging measurement
 776(2008-Jul-31)  38.640   +/- 0.007   Beam injection, poor PLL, Transmitted DC
 782(2008-Aug-02)  38.6455  +/- 0.0012  Beam injection, independent PLL, Transmitted DC
this(2008-Aug-04)  38.64575 +/- 0.00037 Beam injection, independent PLL, Transmitted RF
----------------------------------------------------------------------------------------
Attachment 1: YEND_LAYOUT.png
YEND_LAYOUT.png
Attachment 2: 4th_FSR1.png
4th_FSR1.png
Attachment 3: 4th_FSR2.png
4th_FSR2.png
  786   Sun Aug 3 20:53:54 2008 ranaConfigurationPEMGuralp
We got our repaired Guralp back in the mail from England (s/n T4157). I plugged it in
to Ben's 3-Guralp breakout box (http://www.ligo.caltech.edu/docs/D/D060506-00.pdf) and
verified that it is not oscillating (like it was before) and that it responds to us
jumping around.

The breakout box has way too much gain, however. The ADC wants +/-2 V and the box puts out
~5 Vpk in the night time.

Looking at the schematic, it has a DC gain of 200 and a double whitener (50,50:10,10) so that
there's a gain of 5000 from 50-2000 Hz. The Guralp has a transduction gain of 800 V/(m/s) and
so we can just calculate what the frequency dependent noise figure of the box has to be. I've
pulled it out, put it on the bench, and started reworking it. I am looking for a soldering/
testing volunteer.

The other kink in the problem is that since we want to use this for the adaptive noise cancellation,
we have to make the noise floor of the readout better than the ambient noise by the same factor
with which we want to cancel the noise.
  785   Sat Aug 2 18:37:41 2008 ranaUpdateSUSOSEM Spectra

The attached PDF file is from the .xml files that I found from 7/30. Looks like someone
took some free swinging data and even made nice plots but didn't elog it. Raspberry for you.
The data files are saved in Templates/FreeSwinging/{ETMX,ETMY,etc.}/2008_07_30.xml

The top left plot on the multi-page file all have the same scale so you can see what's happened.
The peaks should all be as measured by Busby in Sep '06
but instead they are as you see here.
Attachment 1: free_080730.pdf
free_080730.pdf free_080730.pdf free_080730.pdf free_080730.pdf free_080730.pdf free_080730.pdf free_080730.pdf free_080730.pdf
  784   Sat Aug 2 16:05:38 2008 ranaConfigurationComputer Scripts / ProgramsmDV update
I did an svn update on our mDV directory. Justin has improved it so that the NDS client binaries
are included for solaris, mac, linux32, and linux64. Now you can just use this version without
having to worry about any path definitions.
  783   Sat Aug 2 13:07:23 2008 KojiConfigurationGeneralThe AP table cleaned
During the construction of the independent PLL I cleaned up some of the unused optics from the AP table. Essentially this should be harmless as they had already been isolated from any beam. They were related to Go's squeezing project and Osamu's MC Transmitted beam measurement.

Nevertherless, if you find any problem on the signals at the AP table (when the ifo returns), I am the person to be blamed.

I am going to update the table layout later next week.
  782   Sat Aug 2 12:53:43 2008 KojiUpdateGeneralAbs. Len. Meas. ~ New PLL at the PSL table
Report of the work last night:
The new heterodyne interferometer on the PSL table was built.
The length of the Yarm cavity was measured with better precision.

-------------
Yarm is locked. The injection beam was aligned. The beat was there at around LT=48.9 [C_deg] of the NPRO.

The new PLL setup on the PSL table has been built. The two beams from the MC incident beam and the injection beam are
mode-matched with lenses. I measured the Rayleigh ranges of the beams by a sensor card and my eyes, and then placed
appropriate lenses so that they can have 5~6 [m] Rayleigh range. They looks a bit too thick but just ok for an inch
optics. The new PLL setup shows ~70% intensity modulation which is enormous. The servo is still SR560-based so far.

Now the PLL has no singular frequency within its range. I could sweep the 4th FSR of the cavity with 500Hz interval. I
was still observing at the transmitted DC.

At each freqency from 15.51MHz to 15.52MHz, a timeseries data of the Yarm transmitted was recorded at sampling of 32Hz for 10
seconds. The figure shows the averaged values of the transmitted DC with errors. An increase of the transmitted power by
3-4% was found. If we consider the resonance is at f_PLL = 15.515 +/- 0.0005 [MHz], this indicates the
arm cavity length of 38.6455 +/- 0.0012 [m].
Y-arm length
e-log    length [m]
-----------------------------
556      38.70   +/- 0.08   Cavity swinging measurement
556      38.67   +/- 0.03   tape & photo
776      38.640  +/- 0.007  Beam injection, poor PLL, Transmitted DC
this     38.6455 +/- 0.0012 Beam injection, independent PLL, Transmitted DC
-----------------------------

NEXT STEPS:
o RF detection at the transmitted
o Better PLL: PLL stability (in-loop / out-of-loop)
o Measurement for the 1st~3rd FSR
o Reproducibility of the measurement
o Higher order mode search
o Check the acuuracy and presicion of the Marconi
Attachment 1: yarm_dc.png
yarm_dc.png
  781   Fri Aug 1 16:33:52 2008 ranaConfigurationPSLPSL Quad change and new script
Here's the sensor ringdown trend from the kick.
Attachment 1: Untitled.png
Untitled.png
ELOG V3.1.3-